# Markov switching zero-inflated space-time multinomial models for comparing multiple infectious diseases

### v1

@@ 4.2 Results
Figure 4 shows fitted values and posterior probabilities of disease presence, $P(S_{kit}=1|\bm{y})$ (see SM Section 5), from three models. SM Section 5 explains how we constructed the fitted values by drawing new space-time random effects and simulating from the fitted model. By comparing panels (a) and (b), the random effects $\phi_{kit}$, from Equations (3)-(4), are an important component of the model and induce a large amount of necessary overdispersion. From panel (c), the model of Zeng et al. (2022), which does not account for correlations in disease presence across space, time and disease, estimates chikungunya as always present in Catumbi, except for the initial state (this is also true for Zika and in all other areas). Therefore, the model of Zeng et al. (2022) is too restrictive to produce the likely complex space-time distribution of disease presence in the city. Comparing panels (a) and (d) shows the MS-ZIARMN model believes the presence of chikungunya was much more persistent compared to Zika, which follows the last two rows of Table 3. One way to think of $P(S_{kit}=1|\bm{y})$ (from the bottom graphs) is as a weight for the observation’s contribution to the likelihood of the multinomial parameters of the model. For example, from panel (a), the MS-ZIARMN model does not use any chikungunya observations before around week 28 in Catumbi to inform the multinomial estimates in Table 2. In contrast, the model of Zeng et al. (2022), from panel (c), will use all observations to inform the multinomial model which likely causes bias due to the large number of 0s, see the discussion at the beginning of Section 4.2.
@@ 5 Discussion
We have proposed a zero-inflated spatio-temporal multinomial model for comparing the transmission dynamics of multiple co-circulating infectious diseases. Our approach can account for long periods of disease absence while investigating how certain factors are related to differences in the transmission intensity of the diseases. Past zero-inflated multinomial models (Tang and Chen, 2019; Zeng et al., 2022; Koslovsky, 2023) have made independence assumptions that are inappropriate for spatio-temporal infectious disease counts, such as assuming categorical presence was independent between observations. We accounted for correlations across space, time and disease, in both the multinomial and zero-inflated components, through a combination of autoregression, regressing on past observations, and assuming the presence of the diseases followed a series of coupled Markov chains. This approach allowed for efficient and computationally feasible Bayesian inference using the FFBS algorithm.
@@
We assume that the cases of the diseases present in an area jointly follow a multinomial model, conditioning on their total. It is more standard in the disease mapping literature to jointly model multiple disease counts with some form of a multivariate Poisson distribution (Jack et al., 2019). There are advantages to both approaches and we give a more detailed comparison of the two (without model fitting) in SM Section 1.2. If one is only interested in comparing the transmission of the diseases, like here, then a multinomial model eliminates many nuisance parameters and any shared space-time factors. A multivariate Poisson approach could estimate the overall effects of a covariate on transmission, not just the relative effects, and does not need to assume one of the diseases is always present. To the best of our knowledge, only Pavani and Moraga (2022) and Rotejanaprasert et al. (2021) have considered zero-inflated spatio-temporal multivariate Poisson models. Both relied on random effects to capture correlations in the data. However, these proposals have some important limitations, including not considering any space-time or space-time-disease interactions and not allowing the probability of disease presence to change over time. These modeling assumptions are inappropriate for our motivating example, and likely others, see Figure 4 for instance. An approach based on random effects could become intractable as one would need to consider random effects correlated across three dimensions, space, time and disease, in both the count and zero-inflated components of the model. Our approach seems much more computationally feasible and could be easily extended to the multivariate Poisson case, see SM Section 1.2. We will focus on these extensions and comparisons with the multinomial model in future work.
@@
In our application, we found that Zika generally had more intense transmission compared to dengue and chikungunya, but was also not able to transmit as well at a lower temperature. This was not a major factor in tropical Rio, see Figure 3, but could be important in North America and Europe. In contrast, chikungunya transmission was relatively higher at lower temperatures meaning it might fare better in colder regions compared to Zika and dengue. Laboratory studies have come to similar conclusions (Tesla et al., 2018; Mercier et al., 2022). Alternative models that did not account for zero inflation or did not model correlations in disease presence (Zeng et al., 2022) did not fit as well and produced estimates inconsistent with our knowledge about the epidemiology of the diseases.
@@
There are also some important limitations to our work. Being an observational study, unmeasured confounding could be an important issue, especially concerning differences in the susceptible populations which are unobservable. We tried using differences in cumulative incidence as a proxy for differences in the susceptible populations, in a sensitivity analysis in SM Section 8, and our results did not change. However, differences in cumulative incidence could be an imperfect proxy due to, for example, changes in reporting rates across space and time. Another limitation is that climate covariates, such as temperature, can have lagged and non-linear effects on the transmission of arboviral diseases (Lowe et al., 2018). As our multinomial model estimates differences in effects on transmission, our approach might be somewhat robust to non-linearity if the non-linear patterns are similar between diseases. It would be challenging to incorporate non-linear and lagged effects into our already complex modeling scheme. Shrinkage methods could help deal with the many parameters that would need to be added to the model by eliminating unimportant effects (Wang et al., 2023).
@@ Acknowledgements
This work is part of the PhD thesis of D. Douwes-Schultz under the supervision of A. M. Schmidt in the Graduate Program of Biostatistics at McGill University, Canada. Douwes-Schultz is grateful for financial support from IVADO and the Canada First Research Excellence Fund/Apogée (PhD Excellence Scholarship 2021-9070375349). Schmidt is grateful for financial support from the Natural Sciences and Engineering Research Council (NSERC) of Canada (Discovery Grant RGPIN-2017-04999). This research was enabled in part by support provided by Calcul Québec (www.calculquebec.ca) and Compute Canada (www.computecanada.ca).

### v2

@@ 4.2 Results
Figure 4 shows fitted values in the top graphs of each panel from three models. SM Section 6 explains how we constructed the fitted values by drawing new space-time random effects and simulating from the fitted model. The bottom graphs of each panel in Figure 4 show the posterior probability the disease was present in the area, $P(S_{kit}=1|\bm{y})$; see SM Section 6. If an area reports cases of disease $k$, then we know that it is present there, that is, $P(S_{kit}=1|\bm{y})=1$ if $y_{kit}>0$. If an area does not report cases, the disease could be absent or present but undetected, the probability of the latter is given by $P(S_{kit}=1|\bm{y})$. By comparing panels (a) and (b), the random effects $\phi_{kit}$, from Equations (3)-(4), are an important component of the model and induce a large amount of necessary overdispersion. From panel (c), the model of Zeng et al. (2022), which does not account for correlations in disease presence, estimates chikungunya as always present in Catumbi, except for the initial state (this is also true for Zika and in all other areas). Therefore, the model presented in Zeng et al. (2022) is too restrictive to capture the likely complex space-time distribution of disease presence in the city. Panel (d) shows Zika likely circulated undetected at the end of 2016. One way to think of $P(S_{kit}=1|\bm{y})$ (from the bottom graphs) is as a weight for the observation’s contribution to the likelihood of the multinomial parameters of the model. For example, from panel (a), the MS-ZIARMN model does not use any chikungunya observations before around week 28 in Catumbi to inform the multinomial estimates in Table 2. In contrast, the model of Zeng et al. (2022), from panel (c), will use all observations to inform the multinomial model, which likely causes bias due to the large number of 0s, see the discussion at the beginning of Section 4.2 and our simulation study in SM Section 4.
@@ 5 Discussion
We have proposed a zero-inflated spatio-temporal multinomial model for comparing the transmission dynamics of multiple co-circulating infectious diseases. Our approach can account for long periods of disease absence while investigating how certain factors are related to differences in the transmission intensity of the diseases. Past zero-inflated multinomial models (Tang and Chen, 2019; Zeng et al., 2022; Koslovsky, 2023) have made independence assumptions that are inappropriate for spatio-temporal infectious disease counts, such as assuming categorical presence was independent between observations. We accounted for correlations across space, time and disease in both the multinomial and zero-inflated components through a combination of autoregression, regressing on past observations, and by assuming the presence of the diseases followed a series of coupled Markov chains. This approach allowed for efficient and computationally feasible Bayesian inference using the FFBS algorithm. We also showed, in a simulation study in SM Section 4, that failing to account for zero-inflation often leads to many important parameters not being covered by the credible intervals, which further lends support to our approach.
@@
We assume that the cases of the diseases present in an area jointly follow a multinomial model, conditioning on their total. It is more standard in the disease mapping literature to jointly model multiple disease counts with some form of a multivariate Poisson distribution (Jack et al., 2019). There are advantages to both approaches and we give a more detailed comparison of the two (without model fitting) in SM Section 1.2. If one is only interested in comparing the transmission of the diseases, like here, then a multinomial model eliminates many nuisance parameters and any shared space-time factors. A multivariate Poisson approach could estimate the overall effects of a covariate on transmission, not just the relative effects, and does not need to assume one of the diseases is always present. As mentioned in the introduction, only Pavani and Moraga (2022) and Rotejanaprasert et al. (2021) have considered zero-inflated spatio-temporal multivariate Poisson models. Both relied on random effects to capture correlations in the data. However, these proposals have some important limitations, including not allowing the probabilities of disease presence to change over time. These modeling assumptions are inappropriate for our motivating example and likely for others as well. For instance, Figure 4 clearly shows the probabilities of disease presence changing over time. An approach based on random effects could become intractable as one would need to consider random effects correlated across three dimensions, space, time and disease, in both the count and zero-inflated components of the model. Our approach seems much more computationally feasible and could be easily extended to the multivariate Poisson case, see SM Section 1.2. We will focus on these extensions and comparisons with the multinomial model in future work.
@@
In our application, we found that Zika generally had more intense transmission compared to dengue and chikungunya, but was also not able to transmit as well at a lower temperature. This was not a major factor in tropical Rio, see Figure 3, but could be important in North America and Europe. In contrast, chikungunya transmission was relatively higher at lower temperatures, meaning it might fare better in colder regions compared to Zika and dengue. Laboratory studies have come to similar conclusions (Tesla et al., 2018; Mercier et al., 2022). Alternative models that did not account for zero inflation or did not model correlations in disease presence (Zeng et al., 2022) did not fit as well and produced estimates inconsistent with our knowledge about the epidemiology of the diseases.
@@
There are also some important limitations of our work. Being an observational study, unmeasured confounding could be an important issue, especially concerning differences in the susceptible populations, which are unobservable. We tried using differences in cumulative incidence as a proxy for differences in the susceptible populations, in a sensitivity analysis in SM Section 9, and our results did not change. However, differences in cumulative incidence could be an imperfect proxy due to, for example, changes in reporting rates across space and time. Another limitation is that weather covariates, such as temperature, can have lagged and non-linear effects on the transmission of arboviral diseases (Lowe et al., 2018). As our multinomial model estimates differences in effects on transmission, our approach might be somewhat robust to non-linearity if the non-linear patterns are similar between diseases. It would be challenging to incorporate non-linear and lagged effects into our already complex modeling scheme. Shrinkage methods could help deal with the many parameters that would need to be added to the model by eliminating unimportant effects (Wang et al., 2023). Also, the effect of weather covariates on the incidence of arboviral diseases may change across space or time (Nobre et al., 2005). To accommodate this, the covariate effects could be modeled as random walks (Nobre et al., 2005) or other spatio-temporal random effects in future work. However, this would be computationally challenging since covariates can enter the model in each of the relative odds and each of the probabilities of presence. Finally, another limitation of our approach is that a multinomial model is not appropriate for predicting future disease counts. To predict future cases of a disease, a multinomial model would need to know the total number of cases across all diseases in the future. A possible solution is to build a separate model for the total count (Schmidt et al., 2022), or a multivariate Poisson model could be used, as discussed above.
@@ Acknowledgements
This work is part of the PhD thesis of D. Douwes-Schultz under the supervision of A. M. Schmidt in the Graduate Program of Biostatistics at McGill University, Canada. Douwes-Schultz is grateful for financial support from IVADO and the Canada First Research Excellence Fund/Apogée (PhD Excellence Scholarship 2021-9070375349). Schmidt is grateful for financial support from the Natural Sciences and Engineering Research Council (NSERC) of Canada (Discovery Grant RGPIN-2017-04999) and from Institut de valorisation des données (IVADO) (PRF-2019-6839748021). L. P. Freitas is supported by the Visiting Researcher Program of the Sergio Arouca National School of Public Health (ENSP), Oswaldo Cruz Foundation (Fiocruz). M. S. Carvalho acknowledges support from the National Council for Scientific and Technological Development (CNPq) - Brazil [307450/2021-0] and Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro (FAPERJ) - Brazil [E-26/203.967/2024]. This research was enabled in part by support provided by Calcul Québec (www.calculquebec.ca) and Compute Canada (www.computecanada.ca).